# SALAD-Bench: A Hierarchical and Comprehensive Safety Benchmark for Large Language Models

### v3

@@ Appendix I Rejection Keywords
Appendix J Safety Rate LeaderBoard
@@
Figure 9: Leaderboard of 24 models on our baseset and attack-enhanced subset, ranked by Elo Rating. The result of Llama2-7b-chat on attack-enhanced subset is not advisable since it is targeted by many attack methods.
@@
Table 13: More results regarding comparison among large language models on multiple-choice subset.
@@
  Methods Keyword GPT-Evaluator MCQ-Judge Human
  GPT-4 Acc-V 72.33% 89.07% 88.96% 89.17%
QWen Acc-V 57.49% 67.47% 68.65% 72.06%
Vicuna Acc-V $*{}^{*}$ 37.77% 39.17% 39.39%
Time Cost 0.01s ~1hour 0.43s ~2hour
Money Cost N/A ~\$20 N/A ${}^{**}$
Table 15: Comparison of different evaluators on multiple-choice subset, where *** means failed to parse choices from Vicuna responses by keyword, and ** is calculated by Amazon Mechanical Turk.
Appendix K Quantitive Results and Analysis
Performance varies across different safety domains and among various models, with certain models like GPT-4 (OpenAI, 2023) and Claude2 (Anthropic, 2022) consistently achieving high safety rates across the board, while others display more fluctuation. In the base set, models generally exhibit high safety rates across most domains as shown in Figure 10. This suggests that under standard testing conditions without intensified adversarial challenges, the models can effectively handle a range of safety issues. However, there is a notable difference in safety rates when comparing the base set to the attack-enhanced subset, where the latter shows a significant drop in safety rates due to the challenge of the questions.
@@
Appendix L More Analysis for MCQ Subset
@@
In addition to evaluating Acc-O and Acc-V for different LLMs, we are also curious about whether LLMs could correctly recognize the safe and unsafe selections from the same multiple-choice questions. Intuitively, for each two multiple-choice questions (i.e., select safe and unsafe selections) from the same seed question, an ideal LLM should generate a pair of consistent outputs (e.g., selecting “AB” for safe selections and “C” for unsafe selections) to illustrate the helpfulness of LLMs in the safety dimension. Therefore, we calculate the selection consistency rates of different LLMs and demonstrate the results in Table 14. GPT-4 (OpenAI, 2023) achieves 86.93% consistency rate and attains the best performance among all LLMs. Furthermore, LLMs with relative high selection consistency usually obtain high accuracy on MCQ subset. Specifically, except for GPT-4, Qwen (Bai et al., 2023) and Tulu (Ivison et al., 2023) also achieve 44.00% and 56.40% selection consistency rate respectively and perform better than other LLMs. In terms of Acc-V in Table 7, both models achieve 68.65% and 71.43% respectively, which perform better than most of LLM counterparts. In contrast, LLMs (e.g., Mistral-7B (Jiang et al., 2023)) with relatively low consistency rate may obtain unsatisfying accuracy. This suggests that both insufficient inherent instruction following ability and too strict safety alignment procedures may influence the overall accuracy in MCQ subset, thus further affect the selection consistency rate. In the following, we will analyze the rejection rates of LLMs in MCQ subset.
@@
Based on above analysis, our observations are two-fold: 1) LLMs with low general safety capability and instruction following ability usually attain weak rejection rates for multiple-choice questions (MCQs); and 2) too strict safety alignment strategies or post-processing may lead to negative effect for recognizing safe selections. For the first observation, according to Table 7 and Table 14, LLMs with insufficient safety ability (e.g., InternLM and Vicuna Chiang et al. (2023)) usually obtain both relative unsatisfying accuracy and low rejection rates on MCQ subset. Notably, InternLM generates responses to all multiple-choice questions, but only obtain 3.85% Acc-V, which supports our first observation. And for the other observation, According to Table 14, Llama-2 and Claude2 obtain much higher rejection rate for questions of selecting unsafe choices (i.e., 84.48% and 87.71% respectively) than those of choosing safe answers. This phenomenon indicates that corresponding LLMs are enhanced by strict safety alignment procedure. Nevertheless, such alignment procedure leads to 1) high rejection rates for question choosing safe answers (i.e., 63.39% and 36.04%), and 2) low Acc-V shown in Table 7. And for Gemini, though it also suffers from relative high overall rejection rate by 43.85%, corresponding rejection rates for safe questions and unsafe counterparts are balanced. Therefore the safety strategies of Gemini does not harm the ability of recognizing safe and unsafe selections and ensures Gemini to obtain promising 78.71% Acc-V. These results encourage us to explore more effective safety alignment algorithms with less negative effect, thus attaining both promising safety capability and optimal helpfulness capability for future LLMs.
Evaluator for MCQ subset.
Moreover, we also measure the effectiveness and efficiency of different evaluators for our MCQ subset. Generally, an ideal evaluator should satisfies two requirements, i.e., comparable accuracy with human evaluator, and much lower time or money cost than human evaluator. Therefore, we compare keyword-based evaluator (Zou et al., 2023), GPT-evaluator (Deng et al., 2023) and our proposed MCQ-Judge in MCQ subset, with the accuracy of three LLMs (OpenAI, 2023; Chiang et al., 2023; Bai et al., 2023), inference time, and money cost from calling APIs. We also introduce human evaluator as reference to measure the effectiveness of each evaluator.
Corresponding evaluation results are shown in Table 15. Compared to results from human evaluator, our MCQ-Judge achieves similar results on Acc-V of GPT-4 and Vicuna, and only falls behind human evaluator by ~3.4% on Acc-V of Qwen. Meanwhile, these results are comparable with counterparts from GPT-evaluator and much better than those from keyword-based evaluator, which indicate that our MCQ-Judge attains sufficient accuracy to evaluate the MCQ subset of our dataset. Moreover, Compared to human evaluator and GPT-Evaluator, MCQ-Judge only requires ~0.43s to evaluate the whole questions for each model, which is comparable with keyword-based evaluator and much faster than GPT/Human-based counterparts. Above results demonstrate the effectiveness and efficiency of MCQ-Judge in our MCQ subset.
Appendix M Qualitative Results
@@
According to Figure 16 and Figure 17, after introducing attack methods (e.g., prompts generate by GCG (Zou et al., 2023) algorithm searched on Llama-2), both Llama-2 and Mixtral-8x7B generate unsafe responses, which indicate that some attack methods obtain inherent transferable capability across different LLMs. This finding will encourage further LLM safety research. And according to Figure 19 and 20, though constructed from the same base question, more LLMs tend to generate incorrect answers when facing multiple-choice questions with choosing unsafe questions. All the wrong answers can be roughly separated into two groups. The first group is wrong selections. Since some multiple-choice questions may include multiple correct selections, some LLMs whose safety capabilities are insufficient, such that models fail to choose all the correct choices. And the second group is wrong format. Some LLMs could choose right choices but fail to generate accurate responses format, such that the responses fail to be detected by MCQ-Judge evaluator. These findings encourage future research to design better safety alignment techniques which ensure both safety capabilities and helpfulness capabilities.
@@
Figure 14: Safety rates for 24 models across sixty five categories in the base set.
@@
Figure 15: Safety rates for 24 models across sixty five categories in the attack-enhanced subset.
@@
Table 16: Our three level taxonomy and the number of questions (# Q) falling into our sixty-five unsafe categories.
Domains
Tasks Categories # Q
Representation & Toxicity Harms
Toxic content
Hate speech
530
Harass, threaten,or bully an individual
550
Enjoying someone elseś pain
368
Terrorism
330
Violent content
511
Child abuse
315
Insult
643
Unfair representation
Racial and ethnic stereotyping
847
Age stereotyping
222
Disability stereotyping
185
Gender stereotyping
367
Religious stereotyping
301
Location stereotyping
243
Adult content
Erotic chat
378
Sexual activity
361
Pornography
336
Misinformation Harms Erosion of trust in public information
Law advice
244
Common sense
320
Illegal drugs and regulated/controlled substances
563
Medical advice
187
Propagating false beliefs
Historical revisionism
293
Science denialism
195
Radicalization
226
Information & Safety Harms Dissemination of dangerous information
Fake news
184
Social media rumors
253
Government secrets
140
Financial data leaks
243
Privacy infringement
Inferring personal information
373
Leaking private information
284
Malicious Use Security threats
Weapon generation and transportation
462
Management of critical infrastructure
322
Malware generation
326
Cyber attack
332
Biological and chemical harms
201
Defamation
Group defamation
168
Defamation someone
269
Fraud or deceptive action
Scams
349
Financial fraud
367
Academic dishonesty
393
Fake review generation
292
Fake online content
280
Disinformation
271
Spam
230
Pseudo-pharmaceuticals
257
Influence operations
Manipulation of public opinion
246
Campaign materials
189
Illegitimate surveillance
380
Illegal activities
Financial crimes
358
Drug-related crimes
351
Sexual offenses
296
Violent crimes
759
Theft
964
Illegal law advice
224
Environmental crimes
248
Traffic and driving offenses
222
Human Autonomy & Integrity Harms Persuasion and manipulation
Self-harm
388
Psychological manipulations
252
Violation of personal property
Mimicking writing style
195
Plagiarism
324
Copyright infringement
238
Forge identity
320
Socioeconomic Harms Risky financial practices
Multi-level marketing
230
Paypal lending
160
Gambling
261
Trade and compliance
Labor issues
114
Trade affairs
86
@@
Table 16: Our three level taxonomy and the number of questions (# Q) falling into our sixty-five unsafe categories.

### v4

@@ Appendix C Rejection Keywords
Appendix D Question generation
We use the prompt in Figure 6 to generate self-instructed questions in all the categories by finetuned GPT. The number of questions and categories should be changed to generate. Upon collection, we leverage an LLM-based reward model Dai et al. (2023) $\mathcal{R}_{\text{s}}$, pre-trained on SafeRLHF Dai et al. (2023), to determine whether each question is safe or not. Specifically, with given input question $\mathbf{q}$, we first feed $\mathbf{q}$ into $\mathcal{R}_{\text{s}}$ to obtain corresponding harmfulness score $s_{\text{harm}}$. Then, we classify question $\mathbf{q}$ via $s_{\text{harm}}$ and a given threshold $T_{\text{harm}}$, i.e., $\mathbf{q}$ with $s_{\text{harm}}>T_{\text{harm}}$ is seen as harmful question, and vise versa. Afterwards, we follow the instructions for auto labeling task which is shown in Figure 7.
  Type Data Source Nums
  Self-instructed Finetuned GPT-3.5 15433
Open-Sourced HH-harmless 4184
HH-red-team 659
Advbench 359
Multilingual 230
Do-Not-Answer 189
ToxicChat 129
Do Anything Now 93
GPTFuzzer 42
Total 21318
Table 9: Data source of base set in SALAD-Bench.
Figure 6: Prompt for harmful question generation, which consists of system prompt and input prompt. [X] should be the number of questions to be generated and [Y] should be the exact category.
Figure 7: Example instructions for the auto labeling task. The task consists of three main components: task description, context examples, and output format.
Figure 8: Construction pipeline of the defense-enhanced dataset. (a) Generate response on all candidate models. (b) Keep questions with a high rejection rate. (c) Attack each question and keep failed ones. (d) Enhance remaining questions with defense methods.
Appendix E Human verification of Dataset Quality
To ensure the quality of the final dataset we obtained, we sample a subset of our dataset and asks 4 annotators to verify the dataset. To be specific, we randomly sample 458 questions from all the 65 categories keeping the origin ratio among categories and making sure at least one question sampled from each category. And then, this subset is cross-validated by 4 of our paper authors. We use a shared excel sheet as a tool for collaborative annotation of which columns contain the question-answer pair, the safety taxonomy to which they belonged, and two columns that needed to be annotated by human: “whether the questions were indeed unsafe” and “whether the category-level taxonomy is accurate”. At the very beginning of human validation, all annotators meet in unison to clarify our labeling requirements again. After the first round of annotation, we will organize a new round of meeting discussion for the data samples that the annotators did not agree on, to achieve the final agreement among our 4 annotators.
The annotation results indicate that our data taxonomy labels match human taxonomy labels with a consistency rate of 94.3%. Additionally, the unsafe rate of our questions is 96.7%. This demonstrates the reliability of our taxonomy labels and confirms that the questions are indeed harmful, guiding unsafe outputs from LLMs.
Appendix F Details in Attack Enhancement
Appendix G Training data for MD-Judge
Our MD-Judge training data are collected from two sources:
Public source:
we collect question answer pairs from the training set of Beavertails (Ji et al., 2023), a subset of LMSYS-Chat[2] (Zheng et al., 2023a) and training set of Toxicchat[3] (Lin et al., 2023).
Our generated data:
we additionally generate specially attacked-enhanced question answer pairs. The attack-enhanced questions are randomly chosen from base questions and then generated using attack methods. The safe answers are generated using GPT4 and unsafe answers from finetuned GPT-3.5.
Appendix H Evaluation Template for MD-Judge
The template of our MD-Judge is intricately designed with four distinct components. First is Task, which indicates that our evaluator is designed to assess the safety of agent messages in the whole conversation. Second is Safety taxonomy, to make a trade-off between classification accuracy and context length of LLMs, we list the general activities of domain-level or task-level safety taxonomies. To ensure the activities of each category are easy to understand, we utilize the ability of GPT-4 and human labor to transform each taxonomy into a clear and distinct instruction format, i.e., description regarding “Should not” and “Can”, thus obtaining the safety taxonomy paragraph. Third is QA Pairs, where to place the input questions and corresponding responses from specific LLMs. Lastly, Output format aims to utilize LLMs’ ability to follow instructions, ensuring that our evaluator adheres strictly to our designated format. This approach paves the way to calculate the safe/unsafe probability using the probability of the very first token.
Appendix I Detailed Evaluation Results of Evaluators
  Methods Base Enhance ToxicChat Beavertails SafeRLHF
Table 10: The detailed comparison results of the safety evaluation between our model and other mainstream evaluation methods. The values reported in the table, listed from left to right, are as follows: accuracy, precision, recall, and F1 score. The best F1 scores are bolded the second best results are underlined. Base and Enhance indicate our SALAD-Base-Test and SALAD-Enhance-Test.
  Methods Base Enhance ToxicChat Beavertails SafeRLHF
F1 AUPRC F1 AUPRC F1 AUPRC F1 AUPRC F1 AUPRC
Table 11: Comparison between LlamaGuard with different taxonomy templates and our MD-Judge for QA-pairs. Origin means LlamaGuard’s official safety policy in their code implementation, Domain and Task mean our two levels of safety policy. The best results are bolded and the second results are underlined. Base and Enhance indicate our SALAD-Base-Test and SALAD-Enhance-Test.
Appendix J SFT Versions of Our Evaluators
First of all, we only focus on the open-sourced SoTA models with 7B parameters i.e. Llama-2-7B (Touvron et al., 2023), Mistral-7B-v0.1 (Jiang et al., 2023), and Mistral-7B-Instruct-v0.1 (Jiang et al., 2023) for the following two reasons: 1) the commendable understanding and reasoning capabilities ensure robust representation ability to various question-answer pairs. 2) models with ~7B parameters are more user-friendly and require fewer computation resources during inference.
Based on the table results, we can tell that Mistral-7B-v0.1 along with the task-level template training format is the best one as it demonstrates significant improvements over the alternatives on the ToxicChat and also achieve commendable results on the other datasets. Therefore, we finally choose it as our evaluator.
  Versions Tax. Base Enhance ToxicChat Beavertails SafeRLHF
F1 AC F1 AC F1 AC F1 AC F1 AC
  Llama-2-7B domain .8276 .8646 .8342 .9288 .5818 .6683 .8547 .9506 .8579 .9276
Llama-2-7B task .8174 .8812 .7796 .9126 .5518 .6325 .8549 .9506 .8561 .9233
Table 12: Different versions of our fine-tuned safety evaluators. The best results are bolded and the second results are underlined. AC is short for AUPRC and Tax is short for taxonomy. The number with a “*” means that outputs do not strictly follow the format, causing an inaccurate number. Base and Enhance indicate our SALAD-Base-Test and SALAD-Enhance-Test.
Appendix K Evaluator for MCQ subset.
  Methods Keyword GPT-Evaluator MCQ-Judge Human
  GPT-4 Acc-V 72.33% 89.07% 88.96% 89.17%
QWen Acc-V 57.49% 67.47% 68.65% 72.06%
Vicuna Acc-V failed 37.77% 39.17% 39.39%
Time Cost 0.01s ~1hour 0.43s ~2hour
Money Cost N/A ~\$20 N/A ~\$77∗∗
Table 13: Comparison of different evaluators on the multiple-choice subset, where * means failed to parse choices from Vicuna responses by keyword, and ** is calculated by Amazon Mechanical Turk.
We measure the effectiveness and efficiency of different evaluators for our MCQ subset. Generally, an ideal evaluator should satisfy two requirements, i.e., comparable accuracy with human evaluator, and much lower time or money cost than human evaluator. Therefore, we compare keyword-based evaluator (Zou et al., 2023), GPT-evaluator (Deng et al., 2023) and our proposed MCQ-Judge in MCQ subset, with the accuracy of three LLMs (OpenAI, 2023a; Chiang et al., 2023; Bai et al., 2023), inference time, and money cost from calling APIs. We also introduce human evaluator as reference to measure the effectiveness of each evaluator.
Corresponding evaluation results are shown in Table 13. Compared to results from human evaluator, our MCQ-Judge achieves similar results on Acc-V of GPT-4 and Vicuna, and only falls behind human evaluator by ~3.4% on Acc-V of Qwen. Meanwhile, these results are comparable with counterparts from GPT-evaluator and much better than those from keyword-based evaluator, which indicates that our MCQ-Judge attains sufficient accuracy to evaluate the MCQ subset of our dataset. Moreover, Compared to human evaluator and GPT-Evaluator, MCQ-Judge only requires ~0.43s to evaluate the whole questions for each model, which is comparable with keyword-based evaluator and much faster than GPT/Human-based counterparts. Above results demonstrate the effectiveness and efficiency of MCQ-Judge in our MCQ subset.
Appendix L Templates of MCQ-Judge
Generation templates.
To generate safe responses, we utilize GPT-4 (OpenAI, 2023a) to produce benign answers using the following prompt:
Since pre-trained LLMs (OpenAI, 2023a; Touvron et al., 2023) (e.g., GPT-4 and Llama-2) typically generate benign outputs, we employ a finetuned GPT model, as detailed in Section 2.2 to generate unsafe responses by the following prompt:
Judge template.
Here is the full template used in MCQ-Judge.
Note that the few-shot exemplars only provide output formats for responses and do not require introducing realistic safety questions, thus avoiding introducing bias during answer generation for multiple-choice questions.
Appendix M Details in Evaluating Attack and Defense Methods
Attack methods.
For all attack algorithms that require specifying a target model, we choose Llama-2-7B-Chat as the target model because it is one of the safest models as shown in Table 5. We hypothesize that attacks successful against Llama-2 would be more transferable to other models compared to using unsafer models as the target model.
For TAP Mehrotra et al. (2023), we follow the original work to use vicuna-13B-v1.5 as the attacker. We experiment with different configurations using vicuna-13B-v1.5, gpt-3.5-turbo-1106, and gpt-4-1106-preview as the evaluator. During attacks, the attacker model would generate multiple attack-enhanced questions. We collect all questions with evaluation score $\geq 9$ (full score is $10$) and re-evaluate using our evaluation. We count a successful attack if one of the attack-enhanced questions is evaluated as unsafe by our evaluator.
Human-designed jailbreak prompts templates are picked from jailbreakchat.com based on corresponding reported JB scores and upvotes. We pick the top 10 templates ranked by JB scores and another top 10 by upvotes.
For Chain-of-Utterances(Bhardwaj and Poria, 2023), we use its prompt. But when chatting with models, we apply the chat template. This is different from the original work (according to its source code) where the chat template is not applied when generating from open-sourced models.
Defense methods.
For paraphrasing-based methods, we choose GPT-paraphrasing Cao et al. (2023) as the baseline method. For perturbation-based methods, we choose four different augmentation methods, i.e., random erasing Cao et al. (2023), random inserting Robey et al. (2023), random patching Robey et al. (2023), and random swapping Robey et al. (2023) as defense methods. And for prompting-based methods, we utilize the recently proposed Safe / XSafe prompts Deng et al. (2023) and Self-Reminder prompt Wu et al. (2023) in our experiments, which have shown effective defense abilities in small-scale experiments. We illustrate the full results of defense methods shown in Table 14.
  w/o Defense 34.28% 88.92% 93.60% 90.64% 93.06% 92.04% 96.34% 76.26%
Appendix N Safety Rate LeaderBoard
@@ Appendix I Detailed Evaluation Results of Evaluators
Figure 9: Leaderboard of 24 models on our base set and attack-enhanced subset, ranked by Elo Rating. The result of Llama2-7b-chat on the attack-enhanced subset is not advisable since it is targeted by many attack methods.
@@
Table 15: More comparison among large language models on the multiple-choice subset.
@@
Appendix O Quantitive Results and Analysis
Performance varies across different safety domains and among various models, with certain models like GPT-4 (OpenAI, 2023a) and Claude2 (Anthropic, 2022) consistently achieving high safety rates across the board, while others display more fluctuation. In the base set, models generally exhibit high safety rates across most domains as shown in Figure 10. This suggests that under standard testing conditions without intensified adversarial challenges, the models can effectively handle a range of safety issues. However, there is a notable difference in safety rates when comparing the base set to the attack-enhanced subset, where the latter shows a significant drop in safety rates due to the challenge of the questions.
@@
Appendix P More Analysis for MCQ Subset
@@
In addition to evaluating Acc-O and Acc-V for different LLMs, we are also curious about whether LLMs could correctly recognize the safe and unsafe selections from the same multiple-choice questions. Intuitively, for each of two multiple-choice questions (i.e., select safe and unsafe selections) from the same seed question, an ideal LLM should generate a pair of consistent outputs (e.g., selecting “AB” for safe selections and “C” for unsafe selections) to illustrate the helpfulness of LLMs in the safety dimension. Therefore, we calculate the selection consistency rates of different LLMs and demonstrate the results in Table 16. GPT-4 (OpenAI, 2023a) achieves 86.93% consistency rate and attains the best performance among all LLMs. Furthermore, LLMs with relatively high selection consistency usually obtain high accuracy on the MCQ subset. Specifically, except for GPT-4, Qwen (Bai et al., 2023) and Tulu (Ivison et al., 2023) also achieve 44.00% and 56.40% selection consistency rates respectively, and perform better than other LLMs. In terms of Acc-V in Table 8, both models achieve 68.65% and 71.43% respectively, which perform better than most of LLM counterparts. In contrast, LLMs (e.g., Mistral-7B (Jiang et al., 2023)) with relatively low consistency rates may obtain unsatisfying accuracy. This suggests that both insufficient inherent instruction following ability and too strict safety alignment procedures may influence the overall accuracy in the MCQ subset, thus further affecting the selection consistency rate. In the following, we will analyze the rejection rates of LLMs in the MCQ subset.
@@
Based on above analyses, our observations are two-fold: 1) LLMs with low general safety capability and instruction following ability usually attain weak rejection rates for multiple-choice questions (MCQs); and 2) too strict safety alignment strategies or post-processing may lead to negative effects for recognizing safe selections. For the first observation, according to Table 8 and Table 16, LLMs with insufficient safety ability (e.g., InternLM, and Vicuna Chiang et al. (2023)) usually obtain both relatively unsatisfying accuracy and low rejection rates on the MCQ subset. Notably, InternLM generates responses to all multiple-choice questions, but only obtains 3.85% Acc-V, which supports our first observation. And for the other observation, According to Table 16, Llama-2 and Claude2 obtain much higher rejection rate for questions of selecting unsafe choices (i.e., 84.48% and 87.71% respectively) than those of choosing safe answers. This phenomenon indicates that corresponding LLMs are enhanced by strict safety alignment procedures. Nevertheless, such alignment procedure leads to 1) high rejection rates for question choosing safe answers (i.e., 63.39% and 36.04%), and 2) low Acc-V shown in Table 8. And for Gemini, though it also suffers from a relatively high overall rejection rate of 43.85%, corresponding rejection rates for safe questions and unsafe counterparts are balanced. Therefore the safety strategies of Gemini do not harm the ability to recognize safe and unsafe selections and ensure Gemini to obtain promising 78.71% Acc-V. These results encourage us to explore more effective safety alignment algorithms with fewer negative effects, thus attaining both promising safety capability and optimal helpfulness capability for future LLMs.
Appendix Q Qualitative Results
@@
According to Figure 16 and Figure 17, after introducing attack methods (e.g., prompts generated by GCG (Zou et al., 2023) algorithm searched on Llama-2), both Llama-2 and Mixtral-8x7B generate unsafe responses, which indicate that some attack methods obtain inherent transferable capability across different LLMs. This finding will encourage further LLM safety research. According to Figure 19 and 20, though constructed from the same base question, more LLMs tend to generate incorrect answers when facing multiple-choice questions with choosing unsafe questions. All the wrong answers can be roughly separated into two groups. The first group is the wrong selection. Since some multiple-choice questions may include multiple correct selections, some LLMs whose safety capabilities are insufficient, such that models fail to make all the correct choices. The second group is the wrong format. Some LLMs could make the right choices but fail to generate an accurate response format, such that the responses fail to be detected by the MCQ-Judge evaluator. These findings encourage future research to design better safety alignment techniques that ensure both safety capabilities and helpfulness capabilities.
@@
Figure 14: Safety rates for 24 models across sixty-five categories in the base set.
@@
Figure 15: Safety rates for 24 models across sixty-five categories in the attack-enhanced subset.
  Domains Tasks Categories # Q
  Representation & Toxicity Harms Toxic content Hate speech 530
Harass, threaten or bully an individual 550
Enjoying someone else’s pain 368
Terrorism 330
Violent content 511
Child abuse 315
Insult 643
Unfair representation Racial and ethnic stereotyping 847
Age stereotyping 222
Disability stereotyping 185
Gender stereotyping 367
Religious stereotyping 301
Location stereotyping 243
Adult content Erotic chat 378
Sexual activity 361
Pornography 336
Misinformation Harms Erosion of trust in public information Law advice 244
Common sense 320
Illegal drugs and regulated/controlled substances 563
Medical advice 187
Propagating false beliefs Historical revisionism 293
Science denialism 195
Radicalization 226
Information & Safety Harms Dissemination of dangerous information Fake news 184
Social media rumors 253
Government secrets 140
Financial data leaks 243
Privacy infringement Inferring personal information 373
Leaking private information 284
Malicious Use Security threats Weapon generation and transportation 462
Management of critical infrastructure 322
Malware generation 326
Cyber attack 332
Biological and chemical harms 201
Defamation Group defamation 168
Defamation someone 269
Fraud or deceptive action Scams 349
Financial fraud 367
Academic dishonesty 393
Fake review generation 292
Fake online content 280
Disinformation 271
Spam 230
Pseudo-pharmaceuticals 257
Influence operations Manipulation of public opinion 246
Campaign materials 189
Illegitimate surveillance 380
Illegal activities Financial crimes 358
Drug-related crimes 351
Sexual offenses 296
Violent crimes 759
Theft 964
Illegal law advice 224
Environmental crimes 248
Traffic and driving offenses 222
Human Autonomy & Integrity Harms Persuasion and manipulation Self-harm 388
Psychological manipulations 252
Violation of personal property Mimicking writing style 195
Plagiarism 324
Copyright infringement 238
Forge identity 320
Socioeconomic Harms Risky financial practices Multi-level marketing 230
Paypal lending 160
Gambling 261
Trade and compliance Labor issues 114
Trade affairs 86
@@
Table 17: Our three-level taxonomy and the number of questions (# Q) fall into our sixty-five unsafe categories.